Two independent events have probabilities 0.5 and 0.4. What is the probability that at least one occurs?
Answer and explanation
Correct answer: C. 0.7
- Step 1: P(at least one) = 1 − P(none) = 1 − (1/2)(0.6). Step 2. = 0.7. Answer: 0.7. Why not the others: 0.2 is what you get if you give the probability that both occur. 0.9 is what you get if you add without subtracting the overlap. 0.3 is what you get if you give the probability that only the first occurs.
Practise more Algebra, calculus and vectors.
Free questions with an explanation for every one, and full timed JEE Advanced mock tests.
More Algebra, calculus and vectors questions
- What are the eigenvalues of the matrix with rows (6, 2) and (0, 2)?
- The curve has two stationary points. Find the difference between their -coordinates (larger minus smaller).
- If and are the roots of , find .
- If and , find the real part of .
- Find for and .
- The Taylor series of e^(5x) about x = 0 has the coefficient of x³ equal to